OpenStudy (anonymous):

Most mornings, malik swims 3/4 mile. today he had only enough time to swim 5/6 of that distance. how far did malik swim

OpenStudy (anonymous):

ok so this is fraction multiplication: you are trying to figure out what 5/6 of 3/4 is so you set it up as: (5*3)/(6*4)=15/24 miles, but we can reduce by dividing both top and bottom by 3: 15/3=5 24/3=8 so the answer is 5/8 miles

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np